OpenStudy (anonymous):

-5 over3 + 3 over 8? show work

OpenStudy (anonymous):

(-5/3)8+(3/8)3

OpenStudy (anonymous):

you know you can always use a calculator

OpenStudy (anonymous):

do it satellite

OpenStudy (anonymous):

do it olglon!! pleaase

OpenStudy (anonymous):

so just to clarify do you wnat this on decimal form or as a fraction?

OpenStudy (anonymous):

fraction

OpenStudy (anonymous):

-5/3 / 3/8 = -5/3 / 8/3= -40/9 now divide -40/9 = 4 -8/9

OpenStudy (anonymous):

whats the answer??

OpenStudy (anonymous):

4 -8/9 four wholes negative eight over nine

OpenStudy (anonymous):

-1.29

OpenStudy (anonymous):

decimal -4.822

OpenStudy (anonymous):

we was adding not multiplying!! olglon

OpenStudy (anonymous):

hahahaha oh damn I think i need glasses now i could have swore that + was a ÷ ok hold ups

OpenStudy (anonymous):

ok -5/3(8) + 3/8(3) = -40/24 + 9/24 = -31/24 = 1 -7/24

OpenStudy (callisto):

-5/3+ 3/8 = ( -5/3)(8/8) +( 3/8 )(3/3) = -40/24 + 9/24 = -31/24 for addition of fraction, the demoninator should be the same, sp find the LCM of them, that is 24. as you multiply something at the bottom, you need to multiply the same number at the top of the fraction so that its value doesn't change. ps: ( -5/3)(8/8) = (-5/3)x1 = -5/3 and ( 3/8 )(3/3) =(3/8)x1 = 3/8

OpenStudy (anonymous):

ANSWER THE LASTEST QUESTION I JUST PUT UP CALLISTO